### v1

# DITING: A Static Analyzer for Identifying Bad Partitioning Issues in TEE Applications

Abstract.
Trusted Execution Environment (TEE) enhances the security of mobile applications and cloud services by isolating sensitive code in the secure world from the non-secure normal world. However, TEE applications are still confronted with vulnerabilities stemming from bad partitioning. Bad partitioning can lead to critical security problems of TEE, such as leaking sensitive data to the normal world or being adversely affected by malicious inputs from the normal world.
To address this, we propose an approach to detect partitioning issues in TEE applications. First, we conducted a survey of TEE vulnerabilities caused by bad partitioning and found that the parameters exchanged between the secure and normal worlds often contain insecure usage with bad partitioning implementation. Second, we developed a tool named DITING that can analyze data-flows of these parameters and identify their violations of security rules we defined to find bad partitioning issues. Different from existing research that only focuses on malicious input to TEE, we assess the partitioning issues more comprehensively through input/output and shared memory. Finally, we created the first benchmark targeting bad partitioning, consisting of 110 test cases. Experiments demonstrate the DITING achieves an F1 score of 0.90 in identifying bad partitioning issues.
Trusted Execution Environment, bad partitioning, static analysis, data-flow, shared memory
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: TOSEMisbn: 978-1-4503-XXXX-X/2018/06ccs: Security and privacy Software and application securityccs: Security and privacy Software security engineeringccs: Software and its engineering Software evolution
1. Introduction
Trusted Execution Environment (TEE) has emerged as an essential component in enhancing the security of mobile applications and cloud services (Huang et al., 2024; Santos et al., 2014; Dong et al., 2023). In a typical TEE application, code is usually divided into secure and non-secure parts. The non-secure code operates in the non-secure environment, also known as the normal world, while the secure code runs within TEE. This mechanism provides significant security advantages by isolating the secure code from the normal world, preventing attackers from accessing sensitive data or disrupting secure functions, such as key management (Luo et al., 2018; Park et al., 2020), cryptography (Volante et al., 2024; Vella et al., 2021), and memory operations (Zhao and Mannan, 2019; Feng et al., 2024). To achieve this isolation, TEE leverages hardware-based features like memory partitioning and access control enforced by the CPU (Zhang et al., 2024; Atamli-Reineh et al., 2017). When the code in the normal world needs to access secure functions or data, it must interact with TEE using specific interfaces. Therefore, partitioning strategies of TEE applications that incorporate accurate isolation (Zhao et al., 2019; Geneiatakis et al., 2012; Kim et al., 2018) and access control (Liu et al., 2023; Geneiatakis et al., 2012) mechanisms can be helpful in reducing the attack surface. For instance, Zhao et al. (Zhao et al., 2019) designed SecTEE, a solution that integrates computing primitives (e.g., integrity measurement, remote attestation, and life cycle management) into TEE. This design enables SecTEE to resist privileged host software attacks from the normal world through isolation. Liu et al. (Liu et al., 2023) proposed a multi-owner access control scheme for access authorization in Intel SGX, which can prevent unauthorized data disclosure. The above research has shown that well-designed partitions of TEE applications can help mitigate risks by establishing clear boundaries and secure communication protocols between the normal world and TEE.
However, TEE applications still face some threats. For example, Khandaker et al. (Khandaker et al., 2020) summarized four attacks that can be implemented in Intel SGX: Concurrency, Order, Inputs, and Nested (COIN attacks). In these attacks, malicious TEE users can exploit the lax detection vulnerabilities of non-secure inputs (e.g., buffer and value with incorrect size) in Intel SGX to interfere with executing programs in TEE. Cerdeira et al. (Cerdeira et al., 2020) conducted a security investigation on some popular systems based on TrustZone. They found that input validation weaknesses and other bugs in TrustZone-assisted systems can be used to obtain permissions of TEE. For example, Huawei TEE allows a TEE application to dump its stack trace to the normal-world memory through debugging channels, leaking sensitive information.
Figure 1. A bad partitioning case of input validation weaknesses.
Through analysis of these cases, we found that vulnerabilities in TEE applications often arise from insecure data communication between the normal world and TEE (Fleischer et al., 2020; Jang and Kang, 2020), which usually relies on three types of parameters: input/output and shared memory (Fan et al., 2020). These parameters are typically managed by the normal world, making them susceptible to unauthorized access. Hence, the usage of these parameters without isolation and access control (e.g., unencrypted data output, input validation weaknesses, and direct usage of shared memory) can introduce bad partitioning issues to TEE applications. For example, there is a bad partitioning case in Fig. 1. The normal-world code validates the size of the input buffer before invoking TEE, but this validation step is missing inside the TEE code. While attackers who gain control over the normal world have the ability to intercept sensitive data or gain unauthorized access to secure operations (Cerdeira et al., 2022), they can tamper with the size of the input buffer before the TEE invocation and disrupt the execution of the TEE application through malicious input. Such scenarios highlight the critical importance of ensuring that partition boundaries are well-defined and communication channels are fortified against potential breaches. Therefore, this paper discusses TEE bad partitioning issues based on two key aspects: improperly ported or unported code into TEE, and the lack of in-TEE code which securely interfaces the normal world and TEE. Unfortunately, there are currently no tools to analyze and identify TEE partitioning issues.
To address these challenges and enhance TEE security, we formulated three rules for detecting these issues. Following these rules, we designed a data-flow analyzer to identify bad partitioning in TEE applications by determining whether input/output and shared memory parameters passed between the normal world and TEE violate the defined rules. This is the first automatic tool that can identify bad partitioning issues in TEE applications. Our contributions of this paper include:
  • The first systematic summary of partitioning issues. With the real-world vulnerable code where input, output, and shared memory parameters are handling in sensitive functions and data of TEE, we summarized that the bad partitioning issues in TEE are often caused by the improper usage of these parameters. Unlike previous research (Khandaker et al., 2020; Cerdeira et al., 2020), which primarily focused on malicious inputs to TEE, our approach addresses the potential vulnerabilities brought by three parameters within TEE, offering a more comprehensive solution to bad partitioning issues.
  • Proposing a parameter data-flow analyzer to identify partitioning issues. Based on parameter types, we established a set of rules that these parameters must adhere to as they flow through sensitive functions and and interact with sensitive data. By applying data-flow analysis to determine whether these parameters violate the established rules caused by partitioning issues, we can effectively identify instances of bad partitioning in TEE applications. We named it DITING, a fantastic beast has the ability to distinguish between good and evil from Chinese Mythology Journey to the West.
  • Creating the benchmark for bad partitioning detection. Due to the lacking of datasets containing TEE partitioning issues, we created the first Partitioning Errors Benchmark named PartitioningE-Bench, which consists of 110 test cases, and 90 cases of them cover 3 types of bad partitioning vulnerabilities. Experimental results demonstrate that DITING achieves an F1 score of 0.90 in identifying bad partitioning issues. Using DITING, we also find 55 issues in the real-world programs and are confirming them with the developers.
  • The source code of DITING and PartitioningE-Bench are available on https://github.com/CharlieMCY/PartitioningE-in-TEE.
The rest of the paper is organized as follows. We first describe the data interaction model and bad partitioning issues of TEE applications in Section 2. Next, Section 3 presents the system design of DITING and Section 4 evaluates the system. Subsequently, we discuss the threats to validity and related works in Section 5 and Section 6, respectively. Finally, we conclude the paper and mention future work in Section 7.
2. Background and Motivation
2.1. TEE Data Interaction
As shown in Fig. 2, the normal world and TEE are two independent environments, separated to ensure the security of sensitive functions and data. In this architecture, the communication between TEE and the normal world involves three types of parameters: input, output, and shared memory (Fan et al., 2020).
Figure 2. Data communication between TEE and the normal world.
Input parameters are used to transfer data from the normal side to TEE, while outputs handle the results or send processed data back. Inputs and outputs are simple mechanisms that allow users to temporarily transfer small amounts of data between the normal world and TEE. However, temporary inputs/outputs realize data transfer through memory copying, slightly lowering the performance of TEE applications due to additional memory copy. They are mainly used to transmit lightweight data, such as user commands and data for cryptographic operations.
Shared memory provides a zero-copy memory block to exchange larger data sets (e.g., multimedia files or bulk data) between two sides (TrustedFirmware.org, 2025). It allows both sides to access the same memory space efficiently, which avoids frequent memory copying. Shared memory can also remain valid across different TEE invocation sessions, making it suitable for scenarios that require data to be reused multiple times.
Moreover, an SDK is responsible for managing these parameters and communication in the normal side. For example, TrustZone-based OP-TEE uses TEEC_InvokeCommand() function to execute the in-TEE code, enabling the shared memory or temporary buffers to transfer data between the two sides (Sebastian et al., 2019). Then, OP-TEE can handle these interactions through its internal APIs, which process incoming requests, perform secure computations, and return responses to the normal world. Similarly, Intel SGX utilizes the ECALL and OCALL interfaces to achieve these functionalities (Regano and Canavese, 2024). ECALLs allow the normal world to securely invoke functions within the SGX enclave, passing data into the trusted environment for processing, while OCALLs enable the enclave to request services or share results with the untrusted normal world.
Table 1 illustrates some differences between input/output parameters and shared memory. It is important to note that, while the normal side cannot directly access the memory copies of input and output parameters in TEE, an attacker with the permissions of the normal world can still tamper with the inputs before they are transmitted to TEE or intercept and read the outputs after they are returned from TEE (Lu et al., 2023; Ma et al., 2024). Additionally, since shared memory relies on address-based data transfer between the two sides, any modifications made to the data on one side will be instantly mirrored on the other opposing side. Therefore, data interactions controlled by the normal world code are the root cause of the bad partitioning issues in TEE applications.
Table 1. Comparison of input/output parameters and shared memory.
Feature Input/Output Shared Memory
Data Size Small Big
Efficiency Low, memory copying required High, no need of additional memory copy
Life-time Temporary, valid for a single TEE invocation Valid for a long time and shared among several TEE invocations
Privilege The normal world cannot reach memory copy in TEE Both the normal world and TEE can access at the same time
2.2. Bad Partitioning in TEE Projects
By analyzing 10 TEE projects from github.com, we identified three primary types of bad partitioning issues: ① unencrypted data output, ② input validation weaknesses, and ③ the direct usage of shared memory. We list some projects and issues they face in Table 2, and discuss each of these issues in detail below. In this paper, we use projects developed for ARM TrustZone as examples. The issues we identify and the proposed solutions work for other platforms as well.
Table 2. Surveys of bad partitioning issues in real-world TEE projects. ● means that the project contains the corresponding issue, and ○ indicates not containing the issue.
Projects Unencrypted Data Output Input Validation Weaknesses Direct Usage of Shared Memory
optee-sdp
optee-fiovb
basicAlg_use
external_rk_tee_user
secvideo_demo
darknetz
acipher
Lenet5_in_OPTEE
hotp
random
2.2.1. Unencrypted Data Output
Since a TEE application can transfer data from TEE to the normal side using output parameters or shared memory, any unencrypted data is at risk of exposure to attackers during the output process, potentially compromising the confidentiality and security of sensitive data.
1char dump[MAX_DUMP_SIZE];
2op.params[0].tmpref.buffer = (void *)dump;
3op.params[0].tmpref.size = MAX_DUMP_SIZE - 1;
4res = TEEC_InvokeCommand(&session, TA_SDP_DUMP_STATUS, &op, &err_origin);
(a) The array dump is an output parameter in normal-world Code.
1// params[0].memref is a memory copy of op.params[0].tmpref in the normal world
[…]
3// params[0].memref will be copied to op.params[0].tmpref after return
4return TEE_SUCCESS;
(b) TEE data delta_refcount is directly copied to the output buffer.
Figure 3. The example of unencrypted data output from optee-sdp.
1char g_AesOutpUT[256] = {0};
2// op.params[1].tmpref.buffer is an input buffer larger than test buffer in TEE
3op.params[1].tmpref.buffer = g_AesOutpUT;
4res = l_CryptoVerifyCa_SendCommand(&op, &session, CMD_AES_OPER);
(a) The input buffer and its size are defined outside TEE.
1char test[] = {0x01, 0x02, 0x03, 0x04, 0x05};
2// params[1].memref.buffer is the copy of op.params[1].tmpref.buffer in the normal world
3TEE_MemMove(params[1].memref.buffer, test, sizeof(test));
(b) TEE_MemMove is called without verifying the size of input buffer.
Figure 4. The example of input validation weaknesses from basicAlg_use.
1unsigned char membuf_input[] = "from_CA_to_TA";
2// definition of shared memory
3TEEC_SharedMemory shm = {
4 .size = sizeof(membuf_input),
5};
6TEEC_AllocateSharedMemory(&ctx, &shm);
7// shared memory can be changed outside TEE
8memcpy(shm.buffer, membuf_input, shm.size);
9op.params[2].memref.parent = &shm;
10res = TEEC_InvokeCommand(&sess, TA_TEST_APP_FILL_MEM_BUF, &op, &err_origin);
(a) Shared memory is managed in the normal world.
1// params[2].memref.buffer is shared memory
2void *buf = params[2].memref.buffer;
3unsigned int sz = params[2].memref.size;
4// params[2].memref.buffer may have been changed outside TEE
5if (!TEE_MemCompare(buf, "from_CA_to_TA", sz)) {
[…]
7}
(b) TEE_MemCompare is invoked without checking the buffer contents.
Figure 5. The example of the direct usage of shared memory from external_rk_tee_user.
Fig. 3 shows a scenario where if the TEE data is copied to an output buffer without being encrypted, then the attacker can directly obtain the plaintext from dump in the normal world. Therefore, implementing proper security measures is essential to prevent unauthorized access or data leakage during the TEE output.
2.2.2. Input Validation Weaknesses
As mentioned in Section 2.1, an attacker with normal-world permissions has the ability to tamper with any data intended for input into TEE. Fig. 4 illustrates that even though the normal-world code defines a much larger buffer than text buffer used in TEE, it is still susceptible to buffer overflow attacks that in-TEE code directly performs memory copy operations in the TEE code without validating the size of the input buffer params[1].memref.buffer. This occurs because the TEE code may incorrectly assume that the input buffer is secure and reliable without explicitly checking its size. If an attacker maliciously alters the input buffer size, memory copy operations within TEE could exceed the boundaries of the inputs, leading to severe consequences such as crashes and data corruption. Similar precautions should be taken when using input values as indices for array access in TEE. Moreover, checking the input value treated as the size argument of the memory request function TEE_Malloc(size) can effectively reduce the probability of memory allocation failure.
To mitigate this risk, it is necessary for the TEE code to validate all inputs before performing any memory operations, regardless of whether input checking is partitioned to the normal side. This essential validation ensures that any tampering or unexpected changes to the input data in the normal world do not compromise the security of TEE.
2.2.3. Direct Usage of Shared Memory
Shared memory, as a zero-copy memory block, provides the same physical memory space accessible to both the secure and non-secure partitions. Since this memory is maintained by the normal-world code, bad partitioning can introduce significant vulnerabilities. If the TEE code directly uses shared memory without validating its contents, an attacker could manipulate the shared memory data and influence the execution of TEE applications.
Fig. 5 shows that variable buf is initialized by a shallow copy of the shared memory. Therefore, the contents of buf can be directly modified from outside TEE. When buf is passed into TEE_MemCompare function, the TEE application may not return the expected results. In order to solve this problem, shared memory should be deep copied to the in-TEE buffer, ensuring that the data used in TEE is isolated from external manipulations.
3. System Design
In this section, we present the design of the proposed DITING to identify bad partitioning issues of TEE projects. With the help of CodeQL (Youn et al., 2023; GitHub, Inc., 2025), a code analysis engine developed by GitHub, we construct data-flows of parameters passed between the normal world and TEE based on Algorithms 2 to 4 in Section 3.3.1. Then, by analyzing the data-flows, DITING can detect insecure operations caused by bad partitioning. The design process involves 6 key steps shown in Fig. 6:
Figure 6. Overview workflow of DITING.
Preparation Step: Rule Establishment. Sensitive functions and data (e.g., array access, memory operations and key mentioned in Fig. 6) are critical to security and privacy. For example, sensitive data should not be directly passed to the normal world through output parameters. Rules are established to define permissible operations on these sensitive functions and data.
Step 1: Parameter Type Identification. The initial step is identifying the parameters exchanged between the normal world and TEE. These parameters are classified into three types based on Section 2.1.
Step 2: Data-flow Generation. Data-flows of input/output and shared memory parameters will be generated. This will trace the path of each parameter and map its interactions with different components.
Step 3: Parameter Analysis. The next step involves analyzing whether the parameters pass through functions or interact with data related to rules in the preparation step.
Step 4: Rule Matching with Parameter Data-flows. If a parameter passes through rule-related components, we match its data-flow against the established rules in the preparation step to identify violations. Parameters that violate the rules will pinpoint bad partitioning issues, such as leakage of sensitive data to the normal world and misuse of shared memory leading to unexpected execution of in-TEE code.
Step 5: Reporting Bad Partitioning Issues. Based on the rule violations detected in the previous step, DITING generates detailed reports highlighting bad partitioning issues.
By following this process, DITING ensures robust detection of partitioning flaws in TEE projects, aiding developers in adhering to security policies and mitigating potential vulnerabilities. In the following, we will describe the implementation of DITING in detail.
3.1. Rule Establishment
Given the discussion on bad partitioning in Section 2.2, we need to define a set of rules to detect these issues. These rules will focus on identifying incorrect parameter interactions with in-TEE sensitive functions and data that arise from bad partitioning.
Rule 1: Encrypting before Output. In this scenario, we mainly consider two critical functions — snprinf() and TEE_MemMove(), and assignment statements. The following code gives an implementation of secure in-TEE data output:
1char key[] = "123456";
2// buf is the ciphertext of key
3TEE_CipherUpdate(key, strlen(key), buf, &size);
4TEE_MemMove(params[0].memref.buffer, buf, size)
In line 4, key is encrypted by the TEE Internal API TEE_CipherUpdate() before copying to an output buffer via TEE_MemMove(dest, src, size), in which output params[0].memref.buffer is the destination buffer dest, and in-TEE data key is the source buffer src. Similarly, for the function snprintf(dest, size, format, args...), all arguments args must be encrypted before they are formatted to dest, which is a TEE output. Meanwhile, when an output value is the left value of an assignment statement, the corresponding right value should also be encrypted data. Rule 1 is shown as follow:
(1) $(e(buf)\Rightarrow w(o,buf))\land(e(v)\Rightarrow assgin(o,v))$
where $buf$ and $v$ are the contents of memory copy and assignment, $e()$ represents the encryption function, $w()$ and $assgin()$ will transfer $buf$ and $v$ to the output parameter $o$. The symbol $\Rightarrow$ defines that the left operation must be executed before the right operation.
Rule 2: Checking Value before in-TEE Use. According to the analysis in Section 2.2.2, even if the input value is checked in the normal world, the attacker still has the ability to tamper with the input and affect the execution of in-TEE code.
1char buf[] = "123456";
2unsigned int size = strlen(buf);
3if (params[0].memref.size < size) {
4 return TEE_ERROR_BAD_PARAMETERS;
5}
6TEE_MemMove(params[0].memref.buffer, buf, size);
7...
8// a special case
9size = params[1].memref.size;
10char *str = TEE_Malloc(size, 0);
11TEE_MemMove(str, params[1].memref.buffer, size);
The above code provides the example of input value check. Before any input data is used, there should be a conditional statement containing that input. The usage of input values should be noted, including operations like indexing for array access, memory copying, and memory allocation. This leads us to define Rule 2:
(2) $isChecked(i)\Rightarrow array(i)\vee m(i)$
Here, $isChecked()$ denotes the input $i$ has been checked in a conditional statement; then, it can be used in array access $array()$ or memory operation functions $m()$, such as TEE_MemMove() and TEE_Malloc(). There is also a special case in lines 10 to 12. An memory space is allocated with the same size as the input buffer. Then, the input will be copied to the allocated memory. In this case, it is secure for an in-TEE dynamically allocated buffer based on the size of input buffer to accept input data, even if no conditional check is performed on the input. Therefore, the issue results need to filter this situation by the condition: $\neg(malloc(i)\Rightarrow m(i))$.
Rule 3: Using Deep Copy of Shared Memory. Since shared memory supports the feature of zero-copy, using a shallow copy of shared memory within TEE allows both the normal world and TEE to access the same physical memory. This creates a direct link between the trusted and untrusted environments, and breaks the isolation of TEE.
1unsigned int size = params[1].memref.size;
2char *buf = TEE_Malloc(size);
3TEE_MemMove(buf, params[1].memref.buffer, size);
4...
5if (!strcmp(buf, "from_CA_to_TA")) {
6 ...
7}
As shown in the above code, the first step after transferring a shared memory buffer into TEE should be creating a deep copy of the shared memory. This ensures that TEE operates on an independent copy of the data, maintaining its isolation and preventing any unauthorized access or influence from the normal world. In Rule 3, an in-TEE buffer $buf$ needs to be initialized via function $deepCopy()$ with shared memory $shm$. Then, any memory operations $m()$ could be performed on $buf$:
(3) $deepCopy(buf,shm)\Rightarrow m(buf)$
3.2. Parameter Type Identification
Since DITING requires constructing and analyzing data-flows for different types of parameters, it needs to first identify the type of each parameter transferred between the normal world and TEE. For example, in TrustZone-based OP-TEE, there are 10 parameter names, and we classify them into 3 types, which are listed in Table 3.
Table 3. Types of parameters in TrustZone-based OP-TEE.
Name Type
TEEC_VALUE_OUTPUT Output
TEEC_MEMREF_OUTPUT
TEEC_VALUE_INPUT Input
TEEC_MEMREF_INPUT
TEEC_VALUE_INOUT Output & Input
TEEC_MEMREF_INOUT
TEEC_MEMREF_WHOLE Shared Memory
TEEC_MEMREF_PARTIAL_INPUT
TEEC_MEMREF_PARTIAL_OUTPUT
TEEC_MEMREF_PARTIAL_INOUT
Input: $F$, all function TEEC_InvokeCommand() calls;
          $S$, all definition statements of parameter type
Output: $T$, types of parameters passed into each function call $f$ of $F$
1 foreach $f\in F$ do
       // $T_{f}$ means the f-th element in $T$, and $S_{0}$ is the first element of $S$
2       $T_{f}\leftarrow S_{0}$;
3       foreach $s\in S$ do
4             if $s$.isBefore($f$) $\land$ $T_{f}$.isBefore($s$) then
5                   $T_{f}\leftarrow s$;
6
7
8
Algorithm 1 Determine the parameter types passed into OP-TEE.
1op.paramTypes = TEEC_PARAM_TYPES(...);
2res = TEEC_InvokeCommand(&session, TA_CMD, &op, &err_origin);
The above code shows that the parameter types are defined by the line 1 statement before the TEE invocation function TEEC_InvokeCommand() in line 2. Since the normal-world code may contain multiple pairs of parameter definition and TEE invocation, for a TEEC_InvokeCommand() call, we need to locate the parameter type definition statement that precedes it and is closest to it. Algorithm 1 gives the determination of the parameter types passed into one invocation. The algorithm iterates through all parameter definitions and TEE invocation statements, identifying the corresponding parameter definition statement for the current TEE invocation according to the condition in line 4, which means there are no other parameter definition statements between a parameter definition and the TEE invocation. Then, we can generate the data-flows of these parameters by code analysis tools, such as CodeQL.
3.3. Detection of Bad Partitioning
3.3.1. Data-flow Generation
In this step, we first build the data-flow of each parameter passed between the normal world and TEE.
Table 4. The sources and sinks of data-flow generation for output parameters.
Sources (1) in-TEE sensitive data
Sinks (1) output buffer, which is the destination buffer for copying the source: $w(o,source)$
(2) output value, which is assigned with the source: $o=source$
(3) all function calls, in which the source is passed: $functionCall(source)$
Input: $F$, all function calls;
          $S$, all assignment statements
Output: $D$, sensitive data list;
            $ODF$, data-flows for outputs
1 $D\leftarrow\emptyset$;
2 $ODF\leftarrow\emptyset$;
3 foreach $f\in F$ do
       // $f$ has the dest argument, which is an output
4       if $f$.hasDestArgument() is output then
             // get the data copied to output
5             $D\leftarrow D\cup\{f.\text{{{geSrcArgument(}}{\emph{}}{{)}}}\}$( ) };
6
7
8foreach $s\in S$ do
       // get the assigned object
9       if $s$.getLeftValue() is output then
             // get the data assigned to output
10             $D\leftarrow D\cup\{s.\text{{{getRightValue(}}{\emph{}}{{)}}}\}$( ) };
11
12
13foreach $d\in D$ do
14       foreach $f\in F$ do
             // check whether the data passed in $f$
15             if $f$.hasArgument($d$) is True then
16                   $ODF_{d}\leftarrow ODF_{d}\cup\{f\}$;
17
18
19
Algorithm 2 Data-flow generation for output parameters.
Table 5. The sources and sinks of data-flow generation for input parameters.
Sources (1) input value
(2) input buffer
Sinks (1) conditional statement with the source: $if(source)$
(2) index of array access, which is the source or expression with the source: $array(source)\vee array(expr(source))$
(3) memory operation function, in which the source passed: $m(source)$
Input: $I$, all input parameters;
          $C$, all conditional statements;
          $A$, all array access statements;
          $F$, all function calls;
Output: $IDF$, data-flows for inputs
1 $IDF\leftarrow\emptyset$;
2 foreach $i\in I$ do
3       foreach $c\in C$ do
             // $i$ is in the controlling exprseeion $e$ of $c$
4             $e\leftarrow c$.getControllingExpr();
5             if $e$.hasOperand($i$) is True then
6                   $IDF_{i}\leftarrow IDF_{i}\cup\{c\}$;
7
8
9      foreach $a\in A$ do
             // $i$ is an index of $a$
10             if $a$.getIndex() is $i$ then
11                   $IDF_{i}\leftarrow IDF_{i}\cup\{a\}$;
12
            // check whether $i$ is in an expression
13             if isInExpr($i$) is True then
                   // get the expression where the $i$ is an operand
14                   $e\leftarrow i$.getExpr();
15                   if $a$.getIndex() is $e$ then
16                         $IDF_{i}\leftarrow IDF_{i}\cup\{a\}$;
17
18
19
20      foreach $f\in F$ do
21             if $f$.hasArgument($i$) is True then
22                   $IDF_{i}\leftarrow IDF_{i}\cup\{f\}$;
23
24
25
Algorithm 3 Data-flow generation for input parameters.
For output parameters, the sources and sinks of their data-flow generation are listed in Table 4. Sensitive data in TEE, such as cryptographic keys and authentication credentials, will be the sources. The sinks consist of output parameters that serve as the targets of source copies and assignment operations. In addition, we also need to obtain what operations the sensitive data in TEE has undergone, that is, which functions they have been called into. Algorithm 2 describes the construction of data-flows for outputs, it has 2 stages:
  1. a)
    Lines 3 to 8 are the procedure of obtaining all in-TEE data that will be copied or assigned to the output parameters.
  2. b)
    Lines 9 to 12 can get all functions that have processed the above data.
Then, we can trace the flow paths of in-TEE data to the outputs $ODF$, and $ODF_{d}$ means the path of $d$.
For input parameters, as illustrated in Table 5, we only need to focus on the statements that they pass through, such as conditions, array access, memory operation function calls, etc. Additionally, the second sink needs to consider two types: a[i] and a[i-1], which are i and the expression composed of i as an array index, respectively. There are three loop blocks at lines 3 to 16 in Algorithm 3. Each of them corresponds to three kinds of sinks in Table 5. Finally, we get $IDF$ which represents the data-flow of each input during execution of in-TEE code, and $IDF_{i}$ denotes data-flow path of $i$.
In the end, the sinks of shared memory data-flows are the buffer initialization statements based on a shallow copy of shared memory, and function calls where they directly treat the shared memory buffer as the argument. According to Table 6, we design Algorithm 4 to generate the data-flows of shared memory parameters. Lines 3 to 5 are responsible for retrieving the status of the first sink, lines 6 to 8 correspond to the second, and $MDF_{m}$ is data-flow of $m$.
Table 6. The sources and sinks of data-flow generation for shared memory.
Sources (1) shared memory buffer
Sinks (1) shallow copy statement with the source: $buf=shm$
(2) all function calls, where directly use the source: $f(shm)$
Input: $M$, all shared memory parameters;
          $S$, all assignment statements;
          $F$, all function calls;
Output: $MDF$, data-flows for shared memory
1 $MDF\leftarrow\emptyset$;
2 foreach $m\in M$ do
3       foreach $s\in S$ do
4             if $s$.getRightValue() is $m$ then
5                   $MDF_{m}\leftarrow MDF_{m}\cup\{s\}$;
6
7
8      foreach $f\in F$ do
9             if $f$.hasArgument($m$) is True then
10                   $MDF_{m}\leftarrow MDF_{m}\cup\{f\}$;
11
12
13
Algorithm 4 Data-flow generation for shared memory.
3.3.2. Rule Matching and Issue Reporting
Once the data-flow of each parameter is constructed, it can be matched against the rules defined in Section 3.1 to identify bad partitioning issues in TEE projects.
Bad partitioning ①. Algorithm 5 describes a way of detecting unencrypted data that will be written to the output. First, for each data-flow path in $ODF$, we initialize a flag $isEncrypted$ to False. This flag tracks whether data in the current path has been encrypted. Then, we iterate through each node $n$ in the current path. If $n$ represents a function call, we need to extract its name and check whether the name contains substrings like enc or cipher. If such a substring is present, set $isEncrypted$ to True. This ensures encryption functions within a data-flow are identified to determine the data has been protected.
For other function nodes, if they have a destination buffer argument that is an output, we should check whether $isEncrypted$ is False. If so, we can record this node in the report $OR$. For $n$, which is an assignment statement, if the left-value of this assignment is an output and $isEncrypted$ of this path is still False while iterating to the node, this node also needs to be put into $OR$. After processing all paths in $ODF$, Algorithm 5 returns the report $OR$, which contains all nodes where unencrypted data is copied to the output.
Input: $ODF$, data-flows for outputs
Output: $OR$, report of unencrypted data output
1 foreach $path\in ODF$ do
2       $isEncrypted\leftarrow False$;
       // $n$ is a node in the $path$
3       foreach $n\in path$ do
4             if isFunction($n$) is True then
                   // get the function name of $n$
5                   if (``$enc"\vee``cipher")$ in $n$.name() then
6                         $isEncrypted\leftarrow True$;
7
8                  if $n$.hasDestArgument() is output then
9                         if $isEncrypted$ is False then
10                               $OR\leftarrow OR\cup\{n\}$;
11
12
13
14            if isAssignment($n$) is True then
15                   if $n$.getLeftValue() is output then
16                         if $isEncrypted$ is False then
17                               $OR\leftarrow OR\cup\{n\}$;
18
19
20
21
22
Algorithm 5 Detection of unencrypted data output.
Bad partitioning ②. Algorithm 6 focuses on detecting potential weaknesses in input validation by checking whether inputs pass through conditional logic before being processed further. It assumes that conditional statements are indicators of input validation and marks nodes of $IDF$ as vulnerable if they precede any such checks.
Similar to Algorithm 5, a flag $isChecked$ is also initialized to False and we retrieve each node in paths of $IDF$. The flag indicates whether input validation has been performed along the path. If the node represents a conditional statement (e.g., an if or switch block), we can set $isChecked$ to True. If $isChecked$ remains False for a node $n$, it suggests that no validation has been performed for this input up to that point. The node $n$ will be added to the report $IR$. In addition, we exclude the special case in Rule 2 through lines 9 to 14.
Input: $IDF$, data-flows for inputs
Output: $IR$, report of input validation weaknesses
1 foreach $path\in IDF$ do
2       $isChecked\leftarrow False$;
3       $isSecure\leftarrow False$;
4       foreach $n\in path$ do
5             if isConditional($n$) is True then
6                   $isChecked\leftarrow True$;
7
8            if $isChecked$ is False then
9                   $IR\leftarrow IR\cup\{n\}$;
10
            // filter the special case
11             if isFunction($n$) is True then
12                   if (``$malloc")$ in $n$.name() then
13                         $isSecure\leftarrow True$;
14
15                  if (``$memove")$ in $n$.name() then
16                         if $isSecure$ is True then
17                               $IR\leftarrow IR-\{n\}$;
18
19
20
21
22
Algorithm 6 Detection of input validation weaknesses.
Bad partitioning ③. Algorithm 7 identifies direct access to shared memory without secure-guards, which can lead to in-TEE data corruption and leakage, or other issues of data integrity and confidentiality. Rule 3 ensures that shared memory is only used after appropriate handling, such as deep copying, to avoid effects from the normal world.
If a node $n$ represents a function call, we will determine whether it performs a deep copy operation where the shared memory $shm$ is the source buffer. If it is true, we could continue to check the next node. For all other cases where $shm$ is directly accessed, then, the node $n$ is tagged on the bad partitioning of directly using shared memory.
Input: $MDF$, data-flows for shared memory
Output: $MR$, report of direct usage of shared memory
1 foreach $path\in MDF$ do
2       foreach $n\in path$ do
3             if isFunction($n$) is True then
                   // $n$ is a deep copy where shared memory is source buffer
4                   if $n$.hasSrcArgument() is $shm$ then
5                         continue;
6
7
8            $MR\leftarrow MR\cup\{n\}$;
9
10
Algorithm 7 Detection of direct usage of shared memory.
Finally, three kinds of bad partitioning within TEE projects are reported by DITING. Moreover, DITING can provide strong scalability; by simply adding new rules to the system, it can be extended to detect a broader range of potential problems. Thence, DITING could adapt to evolving requirements and uncover additional vulnerabilities arisen from bad partitioning, making it valuable for ensuring robust and secure implementations of TEE projects.
4. Evaluations
We have conducted several experiments to prove the effectiveness of DITING. In particular, our evaluation focuses on addressing the following research questions (RQs):
  1. RQ1.
    How effective is DITING?
  2. RQ2.
    How efficient is DITING?
  3. RQ3.
    How DITING performs for the real-world TEE projects?
4.1. Environment Setup
4.1.1. Benchmark for Evaluation
Prior studies (Nong et al., 2022; Afrose et al., 2019) have created benchmarks for various static analysis tools that involve injecting vulnerabilities into code. However, none of the existing studies have created a benchmark for TEE partitioning issues. By referencing the construction methodology of the popular API-misuse dataset CryptoAPI-Bench (Afrose et al., 2019, 2023; Torres et al., 2023), we designed the benchmark PartitioningE-Bench for evaluating the abilities of DITING in bad partitioning detection. We will describe its creation steps below.
1int function(TEE_Param params[4])
2{...
3 params[0].value.a = var;
4 ...
5}
(a) Example code of the in-procedure cases. In this case, the critical data params may be passed to another in-TEE function instead of being used directly after entering TEE, making it more challenging to track its data-flow.
1snprintf(params[1].memref.buffer, params[1].memref.size, "%s-%s-%d", key, vi, var);
(b) Example code of the variadic function case. In this case, snprintf may have multiple arguments that need to be copied to the output buffer. Each argument must be checked to ensure it is encrypted.
1if(params[2].memref.size > size) {
2 memcpy(params[2].memref.buffer, buf, size);
3}
(c) Example code of control-flow-based cases. In this case, the memory operation memcpy is inside a conditional statement that checks the buffer size, so this operation is secure.
Figure 7. Examples of test cases in PartitioningE-Bench.
We divide the test cases into 5 parts, including basic cases, in-procedure cases, variadic function cases, control-flow-based cases, and combined cases. These parts will cover various bad partitioning issues.
  1. (1)
    Basic cases are some simple instances described in in Section 2.2. These vulnerabilities typically occur in statements where parameters are directly used after invoking TEE.
  2. (2)
    In-procedure cases mean that the parameter may be passed into other procedures or methods as an argument, which is shown in Fig. 7(a).
  3. (3)
    Fig. 7(b) gives a variadic function case, in which snprintf() may have more than one argument copied to the output buffer. We need to perform data flow analysis on each of them to find the issues of unencrypted data output.
  4. (4)
    Control-flow-based cases will provide some no-issue cases, which can evaluate the precision of DITING. For example, in Fig. 7(c), line 2 should not be identified as a bad partitioning issues, because it is in an if block, which checks the buffer size.
  5. (5)
    Combined cases will introduce at least two of the above cases.
As illustrated in Table 7, we totally provide 110 cases for PartitioningE-Bench, which are written in C programming language and developed for ARM TrustZone. Among these cases, 90 cases contain bad partitioning issues and 20 cases have no issues.
Table 7. The number of bad partitioning issues and no-issue cases in PartitioningE-Bench.
Bad Partitioning Issues Number of Cases
Unencrypted Data Output 35
Input Validation Weaknesses 29
Direct Usage of Shared Memory 26
No-issue Cases 20
Total Cases 110
4.1.2. Experimental Platform
DITING is implemented by Python in combination with the static analysis tool CodeQL. CodeQL is utilized to generate the data-flow for each parameter based on sources and sinks in Tables 4 to 6, and Algorithms 2 to 4 for data-flow generation. The Python code is responsible for retrieving these data-flows and identifying the locations of bad partitioning issues with the help of Algorithms 5 to 7. All experiments were conducted on a machine running Ubuntu 24.04, equipped with a 48-core 2.3GHz AMD EPYC 7643 processor, and 512GB RAM.
4.2. RQ1: Precision and Recall Evaluation
Table 8. The detection precision and recall of DITING on PartitioningE-Bench. NI is the number of partitioning issues, N is the number of detection results, and TP is the number of true positives. P, R and F1 indicates the precision, recall, and F1 score, respectively*.
Bad Partitioning Issues NI N TP P(%) R(%) F1
Unencrypted Data Output 35 34 33 97.06 94.29 0.96
Input Validation Weaknesses 29 27 23 85.19 79.31 0.82
Direct Usage of Shared Memory 26 30 25 83.33 96.15 0.89
Total 90 91 81 89.01 90 0.90
    Note: $P(\%)=TP/N\times 100$, $R(\%)=TP/NI\times 100$.
Table 8 lists the ground truth count of bad partitioning instances in these TEE projects, alongside the number of issues reported by DITING. It can be seen that DITING achieves the highest precision of 97.06% in detecting bad partitioning of unencrypted data output. However, for the other two issues, more FPs are observed.
Through analysing the examples of misdiagnosed code shown in Fig. 8, we categorize the FPs into the following three types:
1char buf[] = "aabbcc";
2...
3params[0].memref.size = strlen(buf);
4// params[1].value.a is the output, while params[2].value.a is the input
5params[1].value.a = params[2].value.a;
(a) The length of buf and params[2].value.a are not the sensitive data.
1unsigned int size = params[0].memref.size;
2// obtain the third last character in the input buffer
3char c = params[0].memref.buffer[size - 3];
(b) Line 3 is the FP of input validation weaknesses, where size is an input parameter as array index.
1// variable assignment by shared memory value
2unsigned int size = params[3].memref.size;
3// shallow copy of shared memory buffer
4void *buf = params[3].memref.buffer;
(c) Line 2 is the FP of direct usage of shared memory, where params[0].memref.size is a shared memory parameter.
Figure 8. Examples of FPs in DITING reports.
  1. (1)
    Non-sensitive data is assigned to the output: Some data can be assigned to the output without encryption, and this will not compromise the confidentiality of TEE. For example, in Fig. 8(a), it is acceptable that we pass the length of the output buffer to the outside or copy the input directly to the output.
  2. (2)
    Indexing for Input Buffer with Its Size: As shown in Fig. 8(b), although the index of array access is influenced by the input value params[0].memref.size, TEE can ensure that this input will not lead to buffer overflow. This is because that the input params[0].memref.buffer in TEE is actually a memory copy of the normal-side data, which is allocated with a size based on the above input value, and this input value is usually the size of the input data outside TEE. It means that the size of params[0].memref.buffer must not be smaller than the input value params[0].memref.size. Therefore, it is secure to directly use params[0].memref.size as an array index, even if no checks are performed on it which violates Rule 2.
  3. (3)
    Variable Assignment Using Shared Memory: Similar to the input buffer, shared memory also transmits memory address and memory size, from the normal side to TEE by params[3].memref. According to Algorithm 1, both of these parameters are treated as the type of shared memory. Meanwhile, any assignment statement containing shared memory parameters is not allowed in Rule 3. However, the variable assignment in line 1 of Fig. 8(c) can be acceptable, as it differs from the insecure shallow copy of shared memory in line 2.
According to the Recall rate in Table 8, DITING can give more comprehensive coverage of bad partitioning related to the direct use of shared memory than other issues. We also give two examples of FNs in Fig. 9 and discuss them as follow:
1void test(char *dest, char *src)
2{
3 TEE_MemMove(dest, src, strlen(src));
4}
5
6void TA_InvokeCommandEntryPoint()
7{
8 char key[] = "123456";
9 ...
10 char *str = TEE_Malloc(strlen(key) + 1, 0);
11 // case 1
12 test(str, key);
13 ...
14 // case 2
15 test(params[1].memref.buffer, key);
16}
(a) Line 15 is the FN of unencrypted data output, where params[1].memref.buffer is an output parameter.
1char *str[1024] = {0};
2...
3for (int i = 0; i < params[2].memref.size; i++) {
4 str[i] = params[2].memref.buffer[i];
5}
(b) Line 4 is the FN of input validation weaknesses, where i is affected by the input params[2].memref.size.
Figure 9. Examples of FNs in DITING reports.
  1. (1)
    Using Memory Operation Functions In Procedure: The destination buffer dest of the function TEE_MemMove in line 3 of Fig. 9(a) is the formal parameter of function test. Therefore, DITING recognizes the type of dest as a character array. However, as illustrated in lines 10 to 13, there are two instances of using test, one involves copying data to the in-TEE buffer str, while the other will copy data to an output. Since DITING does not identify the actual parameter type of dest as an output, it also fails to report the above second instance as bad partitioning, which violates Rule 1.
  2. (2)
    Array Access Affected by Input: Lines 3 to 5 give a for loop block, which is responsible for copying an buffer byte by byte. At this point, the input value params[2].memref.size does not directly serve as an index for array access but is instead used in the condition of for statement to constrain the value of i. Therefore, Algorithm 3 does not recognize this type of array access statement, that has broken Rule 2.
In summary, if we formulate additional filters for the reported issues in the future, it is possible to effectively reduce FPs that deviate from the predefined rules but do not actually introduce security risks. What’s more, we can refine the rules by incorporating the semantic context of the code, and FNs can be further reduced.
Answer to RQ1: (1) DITING provides the highest F1 score of 0.96 in the detection of unencrypted data output (both precision and recall exceed 90%). (2) The detection of direct usage of shared memory achieves the best recall (above 95%), but with relatively low precision (below 85%). (3) The detection of input validation weaknesses has an F1 score of 0.82. While its precision is above 85%, the recall remains slightly lower at approximately 79%.
4.3. RQ2: Efficiency Evaluation
Figure 10. Time cost of data-flow construction and analysis under different scales of LoC in DITING.
On our benchmark tests, we recorded the time taken by DITING to generate data-flows for three types of parameters and match these data-flows against the corresponding rules.
Algorithms 2 to 4, which are responsible for constructing data-flows for parameters, involve $m$ parameters that need to be tracked and $n$ sensitive functions and data in TEE, resulting in the algorithmic complexity of them being $O(mn)$. As shown in Fig. 10, the time consumed by constructing data-flows increases from 21.53 seconds to 34.95 seconds as the lines of code (LoC) grow. Then, DITING focuses only on critical functions and data during data-flow analysis, such as TEE_MemMove and array access operations. The time spent on data-flow analysis remains relatively low and does not grow significantly with the changes of LoC (from 3.96 milliseconds to 7.13 milliseconds).
Answer to RQ2: As the code scale increases from 322 LoC to 5162 LoC, the time cost of the data-flow construction for bad partitioning analysis increases from 21.53 seconds to 34.95 seconds. However, this time consumption also depends on the performance of the underlying code analysis tool. DITING still exhibits high efficiency for detecting bad partitioning issues in TEE projects (around 5.56 milliseconds).
4.4. RQ3: In-the-Wild Study
We searched for TEE-related projects on GitHub using API keywords associated with TEE. After excluding TEE SDK projects, we selected the top 16 projects with the highest star counts to evaluate the performance of DITING on real-world programs, as shown in Table 9. These projects include 10 projects that have been manually analyzed in Section 2.2, along with 6 additional projects.
In these projects, DITING reports 68 issues caused by bad partitioning and 55 of them are confirmed by ourselves. For example, in optee-sdp project, the log functionality implemented via snprintf copies a large amount of register data into the output parameters, which are accessible only to TEE. This compromises the confidentiality of TEE. Additionally, in projects such as Lenet5_in_OPTEE, external_rk_tee_user, there are instances where input buffers are directly used without validating their length. Attackers could exploit this bad partitioning to cause buffer overflows in TEE applications. What’s more, in secvideo project, encryption and decryption operations are directly performed on video data in shared memory, undermining the isolation provided by TEE.
Table 9. The performance of DITING in detecting three kinds of bad partitioning issues on real-world TEE projects.
Projects LoC Unencrypted Data Output Input Validation Weaknesses Direct Usage of Shared Memory
Report Number True Positive Report Number True Positive Report Number True Positive
optee-sdp 830 11 11 7 7
optee-fiovb 776 2 1 2 2
basicAlg_use 5,162 2 2 1 1
external_rk_tee_user 716 4 2 2 2 1 1
secvideo_demo 645 7 4
darknetz 41,608 1 1
acipher 338 1 0
Lenet5_in_OPTEE 1,610 1 1 3 2
hotp 428 1 1
random 322 1 1
read_key 404 2 1
save_key 529 2 1
socket-benchmark 1,209 5 4
socket-throughput 1,100 5 5
tcp_server 750 2 2
threaded-socket 1,059 5 3
Answer to RQ3: DITING is capable of identifying bad partitioning issues in real-world projects. Through a review of 68 issues reported by DITING in the wild, we have confirmed that 55 of them were true alarms.
5. Threats to Validity
5.1. Internal Validity
We use the PartitioningE-Bench dataset, which includes manually injected vulnerabilities, to evaluate DITING. This could lead to a discrepancy between the distribution of bad partitioning issues in the dataset and those in real-world TEE projects, thereby affecting the evaluation of detection accuracy. It is possible that we incorporate more vulnerabilities from real-world TEE projects into PartitioningE-Bench to reduce biases introduced by manual injection.
The performance of DITING also heavily relies on the underlying static analysis tools. However, most tools (e.g., CodeQL and Frame-C (Correnson et al., 2024)) typically necessitate successful compilation and expert customization. This dependency limits their applicability to analyzing incomplete or uncompilable code. If these tools fail to accurately generate data-flows, the detection capabilities of DITING could be compromised. LLMDFA (Wang et al., 2024) proposed by Wang et al.makes up for this defect. Combined with the large language model, LLMDFA effectively improves the accuracy of identifying data-flow paths.
5.2. External Validity
The evaluations focus on TEEs based on ARM TrustZone and Intel SGX architectures. The detection performance of DITING on other TEE architectures, such as AMD SEV (Kaplan, 2023; Mofrad et al., 2018) or RISC-V-based TEE (Nasahl et al., 2021; Suzaki et al., 2020), needs further evaluation. In addition, the evaluated projects are relatively small in scale and do not cover complex industrial-level projects. Future studies should evaluate DITING on a broader range of TEE architectures and large-scale real-world projects, thereby improving the external applicability of the experimental results.
5.3. Construct Validity
Our study defines three types of rules (encrypting before output, checking value before in-TEE use, and using deep copy of shared memory) to identify bad partitioning issues. However, in practical development, there may be certain security assumptions that violate these rules without actually compromising security, potentially leading to false positives. For example, directly outputting a static string (e.g., a software version number, user id, etc.) that does not contain any sensitive data is considered secure and does not require encryption. Therefore, if these definitions are not accurate or comprehensive enough, it may lead to biased detection results. We need to update the rules based on the newly discovered bad partitioning issues in actual TEE applications to expand the detection capabilities of DITING.
6. Related Work
6.1. Vulnerabilities in TEE
TEE has become integral to security-critical applications by providing isolated execution environments to protect sensitive data and computations. However, TEE applications are still affected by some vulnerabilities, as demonstrated by extensive research across various TEE technologies. Cerdeira et al. (Cerdeira et al., 2020) conducted a study on vulnerabilities in TrustZone-based systems. They have revealed critical implementation bugs, such as buffer overflows and inadequate memory protection mechanisms. Their work also underscores the role of hardware-level issues like microarchitectural side-channels in compromising TEE security. Fei et al. (Fei et al., 2021) provided a taxonomy of secure risks in Intel SGX and discussed attack vectors, such as hardware side-channel exploits and software code vulnerabilities. Similarly, Khandaker et al. (Khandaker et al., 2020) introduced the concept of COIN attacks against SGX, and they expose the insecurity of untrusted interfaces in TEE projects. Bulck et al. (Van Bulck et al., 2019) assessed the vulnerabilities at the level of the Application Binary Interface (ABI) and the Application Programming Interface (API) that can lead to memory and side-channel attacks in TEE runtime libraries.
To mitigate the risk of attacks on TEE projects, some research has designed tools to detect vulnerabilities within TEE. Briongos et al. (Briongos et al., 2023) present CloneBuster for detecting cloning attacks on Intel SGX applications. CloneBuster can find whether there are same TEE application binaries running on one platform, which may roll back the TEE applications to the previous state. Ghaniyoun et al. (Ghaniyoun et al., 2023) designed TEESec, a pre-silicon framework for discovering microarchitectural vulnerabilities in TEE. TEESec can test underlying microarchitecture against data and metadata leakage across isolation boundaries.
However, most of these studies focus on the security of TEE hardware architecture and development kits that may have side-channel attack vulnerabilities. Research addressing bad partitioning issues in TEE projects remains limited, despite it being an equally critical concern. Poorly designed TEE code partitioning can result in severe security risks, such as the leakage of sensitive data or the injection of untrusted data, which could influence the integrity of TEE application execution. Our work addresses this gap by comprehensively detecting security vulnerabilities caused by bad partitioning in TEE code, particularly in scenarios involving data interactions between the normal world and TEE with input, output, and shared memory.
6.2. TEE Application Partitioning
Some research has attempted to mitigate vulnerabilities arising from bad partitioning by porting entire programs into TEE. For example, CryptSQLite (Wang et al., 2020) is a database system that places the entire database engine into TEE. This approach prevents attackers from stealing or tampering with critical data computed in TEE from the normal world, effectively protecting the confidentiality and integrity of data during database retrieval. RT-TEE (Wang et al., 2022) is also a cyber-physical system that puts the whole flight control code into TrustZone. This can protect the stability of equipment like drones. However, this way significantly increases Trusted Computing Base (TCB) of the applications, contradicting the principle of maintaining a small TCB for trusted code (Shinde et al., 2017; Li et al., 2019).
To address this, some other research made efforts to identify sensitive data and functions that require protection and place them inside TEE, ensuring a balance between security and efficiency. Rubinov et al. (Rubinov et al., 2016) developed an automatic approach for partitioning Android applications into critical code running in TEE and client code in the normal world. They also reduce the overhead due to transitions between the two worlds by optimizing the granularity of TEE code. Sun et al. (Sun et al., 2024) proposed dTEE that enables developers to declare tiered-sensitive variables and functions of existing IoT applications. What’s more, dTEE can automatically transform device drivers into trusted ones. Moreover, DarkneTZ (Mo et al., 2020) is a framework that limits the attack surface of the model against Deep Neural Networks with the help of Trusted Execution Environment in an edge device. They partitioned the model into sensitive layers inside TEE, and other layers executed in the normal world. This will provide reliable model privacy and defend against membership inference attacks.
The above TEE program partitioning schemes ensure software security while also maintaining the efficiency of program execution. However, as discussed in Section 2.2, improper TEE partitioning can still result in vulnerabilities. Therefore, a tool capable of analyzing and identifying bad partitioning issues in TEE implementations is essential.
7. Conclusion and Future Work
Bad partitioning in TEE projects can lead to vulnerabilities during data transmission between the normal world and TEE. To address this problem, we designed DITING, a tool that can identify security risks arising from bad partitioning by analyzing three types of data interactions: output, input, and shared memory. Additionally, we created the first partitioning errors benchmark, PartitioningE-Bench, consisting of 110 test cases to evaluate the effectiveness of DITING. From the evaluation results, DITING can help to find bad partitioning issues in TEE applications. In the future, we aim to explore automated solutions for fixing bad partitioning issues in TEE applications based on the detection results.
Acknowledgements.
This research is supported by the National Research Foundation, Singapore, and the Cyber Security Agency of Singapore under its National Cybersecurity R&D Programme (Proposal ID: NCR25-DeSCEmT-SMU). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of the National Research Foundation, Singapore, and the Cyber Security Agency of Singapore.